OpenStudy (anonymous):

Evaluate the following function at the indicated values. Round to the nearest ten-thousandth. P(t)=e^-(2/3)t (a) P(-3) = (b) P(7) = (c) P(1/2) =

OpenStudy (amistre64):

\[e^x=1+x+\frac{x^{2}}{2!}+\frac{x^{3}}{3!}+\frac{x^{4}}{4!}+\frac{x^{5}}{5!}+...\] might be useful :)

OpenStudy (anonymous):

?

OpenStudy (amistre64):

seeing how e^x is a rather hard thing to compute as is; we can turn it into a polynomial of approximate value.

OpenStudy (amistre64):

but that is neither here nor there

OpenStudy (amistre64):

-2/3 * -3 = 2 e^2 = 1+2+4/2+8/6+16/24 .... etc

OpenStudy (anonymous):

okay so step by step e^-(2/3)-3 -2/3 * -3 2 e^2=7.389

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

as long as the rounding is good ;)

OpenStudy (anonymous):

sweet thank you

OpenStudy (amistre64):

and of course the others carry along the same concept ...

OpenStudy (anonymous):

yes thank you